Hi I need help with Rational Equations. It says to solve the equation choosing a method.
39. 1/4+4/x=1/x
40. -3x/(x+1)=-2/(x-1)
41. x/6-1/x=1/6
Im using Algebra 1 Concept and Skills By McDougal Chapter 11.7. If you can do these problems and explain them to me I'd be very grateful. Thank You!

Answer by tutorcecilia(2152)

With rational expressions, try to get rid of the denominator by finding the LCD.
39. 1/4+4/x=1/x [LCD = 4x]
4x(1/4+4/x=1/x) [multiply each term by the LCD and cancel wherever possible]
4x/4+(4)(4x)/x=4x/x
x+16=4
x+16-4=4-4
x+12=0
[solve for x]
x=-12
.
Check by plugging x=-12 back into the original equation:
1/4+4/(-12)=1/(-12)
4/-48=-1/12
-1/12=-1/12 [checks out]
.
.
.
40. -3x/(x+1)=-2/(x-1)
LCD=(x+1)(x-1)
(x+1)(x-1)[-3x/(x+1)=-2/(x-1)]
(-3x)(x-1)=(-2)(x+1) [simplify]
-3x^2+3x=-2x-2
-3x^2+3x+2x+2=0
-3x^2+5x+2=0
Use the quadratic formula to solve for zero:
x= 2 or x=-1/3
.

41. x/6-1/x=1/6
LCD=6x
6x(x/6-1/x=1/6 )
(6x)(x)/6-(6x)/x=(6x)/6
x^2-6=x
x^2-x-6=0
Use the quadratic formula to solve for x:
x=3 or x=-2